# CVVNet: A Cross-Vertical-View Network for Gait Recognition

### v1

@@ Abstract
Gait recognition enables contact-free, long-range person identification that is robust to clothing variations and non-cooperative scenarios. While existing methods perform well in controlled indoor environments, they struggle with cross-vertical view scenarios, where surveillance angles vary significantly in elevation. Our experiments show up to 60% accuracy degradation in low-to-high vertical view settings due to severe deformations and self-occlusions of key anatomical features. Current CNN and self-attention-based methods fail to effectively handle these challenges, due to their reliance on single-scale convolutions or simplistic attention mechanisms that lack effective multi-frequency feature integration. To tackle this challenge, we propose CVVNet (Cross-Vertical-View Network), a frequency aggregation architecture specifically designed for robust cross-vertical-view gait recognition. CVVNet employs a High-Low Frequency Extraction module (HLFE) that adopts parallel multi-scale convolution/max-pooling path and self-attention path as high- and low-frequency mixers for effective multi-frequency feature extraction from input silhouettes. We also introduce the Dynamic Gated Aggregation (DGA) mechanism to adaptively adjust the fusion ratio of high- and low-frequency features. The integration of our core Multi-Scale Attention Gated Aggregation (MSAGA) module, HLFE and DGA enables CVVNet to effectively handle distortions from view changes, significantly improving the recognition robustness across different vertical views. Experimental results show that our CVVNet achieves state-of-the-art performance, with $8.6\%$ improvement on DroneGait and $2\%$ on Gait3D compared with the best existing methods.
@@ 1 Introduction
In real-world surveillance scenarios, cameras are generally mounted at elevated positions [18], creating complex viewing conditions. As illustrated in Fig. 1, during normal walking sequences, subjects experience continuous distance variations (far-to-near) and vertical angle change (low-to-high) with respect to the camera. This dynamic perspective induces significant deformation in critical anatomical components (e.g., foot trajectories and joint angles), while key body regions (e.g., torso and feet) suffer from severe self-occlusion under increased vertical angles. Our experimental results demonstrate that under these challenging conditions, both skeletal point-based and local texture-based feature extraction approaches exhibit substantially degraded performance. This perspective sensitivity constitutes a fundamental limitation that severely constrains the practical applicability of existing algorithms, presenting significant challenges to recognition robustness in pervasive cross-vertical viewing-angle scenarios in real-world applications.
@@
Existing methods encounter two fundamental limitations under cross-vertical-view scenarios. First, convolutional networks with fixed receptive fields predominantly focus on local details, resulting in insufficient capacity for modeling global structural information. This shortage makes it challenging to maintain spatial consistency of contours under dynamic viewpoint changes, ultimately amplifying self-occlusion effects. Second, although self-attention mechanisms offer superior low-pass filtering characteristics that effectively capture low-frequency global feature, it suffers from an inherent constraint on preserving high-frequency motion details [37, 38]. Such fragmented frequency representation creates a critical dilemma under continuous vertical viewpoint variations: convolutional architectures fail to compensate for cross-view deformations due to their weak sensitivity to global structural information, whereas self-attention mechanisms sacrifice fine-grained motion cues that are essential for discriminative identity features. The combined effect of these limitations substantially undermines the recognition robustness in complex outdoor surveillance environments, where vertical angle variations are prevalent and unavoidable.
@@
To address the aforementioned challenges, we propose a Cross-Vertical-View Network (CVVNet) to address the generalization bottleneck of existing methods in cross-view scenarios by constructing dual mechanisms of frequency-decoupled feature separation and spatially adaptive gated fusion. The key component of CVVNet is the Multi-Scale Attention Gated Aggregation (MSAGA) module which consists of a High-Low-Frequency feature extraction (HLFE) module and a Dynamic Gated Aggregation (DGA) mechanism. In our HLFE module, the low-frequency branch employs spatial compression attention to disentangle the global structural components of feature contours, while the high-frequency branch utilizes depthwise separable convolution to extract local dynamic components. Furthermore, we use DGA to generate frequency-sensitive fusion coefficients through joint spatial-channel modeling, thereby achieving context-aware recalibration of multi-frequency features. Our main contributions are summarized as follows.
@@
  • We design a Cross-Vertical-View Network (CVVNet) to achieve high-low frequency feature representations for both cross-vertical-view and outdoor recognition scenarios.
@@
  • We propose Multi-scale Attention Gated Aggregation module (MSAGA) to extract high-low frequency information from gait features and dynamically modulate the fusion ratios of different frequency components.
@@ 3.1 Overview
The framework of the proposed method is illustrated in Fig. 2. We first employ a 3D convolutional layer to extract shallow features. Subsequently, the Multi-Scale Attention Gated Aggregation module (MSAGA) is utilized to extract and aggregate both high- and low-frequency information from gait features. The processed features are then further refined through P3D blocks [8] to enhance their representations. Following this, a temporal max pooling operation is applied to aggregate the feature map sequence across the temporal dimension, outputting a holistic representation of the input gait sequence.
@@ 3.2 Frequency Feature Extraction based on HLFE
As shown in Fig. 3, the High-Low Frequency feature Extraction (HLFE) module consists of two branches for high frequency and low frequency feature extraction.
@@ 4.1 Datasets and Implementation Details
Implementation Details. Our network implementation and reproduction experiments of state-of-the-art models are both deployed through the OpenGait codebase [7]. We first integrate the DroneGait dataset into this framework, following the preprocessing methodology described in [4]. The input silhouettes are aligned and resized to $64\times 44$ pixels, and then organized into sequences of length $30$ as model inputs. Both the loss functions and data augmentation strategies of CVVNet adhere to the configurations established in GaitBase [7]. To ensure fair training, we set the batch size for all models to 8×16 for DroneGait and 32×4 for Gait3D, while keeping other settings consistent with the original reproduced models.
@@ 4.2 Results and Analysis
Evaluation on Gait3D. The results on the Gait3D dataset further validate our hypothesis: enhancing viewpoint generalization directly translates to improved accuracy in complex outdoor environments. As shown in Tab 3, CVVNet achieves state-of-the-art performance across all metrics, attaining 75.6% Rank-1, 89.1% Rank-5, and 68.3% mAP. This represents a 1.2% Rank-1 improvement over the previous best method DeepGaitV2 (74.4%) and a 2.5% mAP gain, demonstrating particularly strong advantages in handling real-world variations like occlusions and viewpoint shifts.
@@
Cross-domain Evaluation. Our bidirectional cross-dataset evaluation between Gait3D and DroneGait validates CVVNet’s superior domain generalization. When trained on DroneGait-Low and tested on Gait3D, CVVNet achieves a state-of-the-art 20.9% Rank-1 accuracy, surpassing DeepGaitV2’s previous best result (20.5%) by 0.4% while attaining 14.2% mAP—a 0.3% improvement over DeepGaitV2. Notably, it outperforms QAGait by 5.4% absolute Rank-1 accuracy, confirming effective transfer of drone-learned viewpoint invariance to outdoor scenes. In the reverse evaluation trained on Gait3D and tested on DroneGait-High, CVVNet demonstrates consistent superiority: 45.9% NM, 48.4% BG, and 47.5% CL accuracy under high-vertical-view conditions, exceeding DeepGaitv2 by 0.7-0.8% across metrics. For low-vertical-view scenarios, it reaches 98.4% NM and 99.5% BG accuracy, outperforming GaitBase’s prior records by 1.2%. This bidirectional dominance stems from our dual-branch architecture’s adaptive feature disentanglement, effectively bridging drone-ground domain gaps.
@@ 4.4 Visualization
We draw the heatmap to highlight the viewpoint-specific feature activation in Fig. 5. CVVNet dynamically prioritizes lower-body regions at low vertical angles and shifts attention to upper-body zones at high angles. Specifically, in low vertical views, the network focuses on dynamic details of the legs and feet, capturing crucial gait patterns that distinguish individual walking styles. In contrast, under high vertical views, the attention shifts to the head and torso, where distinctive features are more pronounced. This ability to adaptively adjust feature extraction priorities ensures that CVVNet effectively integrates multi-scale features, enhancing its robustness across varying perspectives and its generalization in complex environments. The spectral analysis is shown in Fig. 6, which reveals distinct frequency characteristics: conventional convolution-based methods (GaitGL, GaitBase, DeepGaitV2) exhibit limited frequency diversity with concentrated central energy. In contrast, our CVVNet’s MSAGA architecture demonstrates superior multi-scale frequency extraction, capturing both low- and high-frequency components. Increasing MSAGA layers progressively enriches the the spectral content, enabling more comprehensive gait representations.
Figure 6: Frequency Spectrum Analysis of Different Gait Feature Extraction Methods.

### v2

@@ Abstract
Gait recognition enables contact-free, long-range person identification that is robust to clothing variations and non-cooperative scenarios. While existing methods perform well in controlled environments, they struggle with cross-vertical view scenarios where elevation changes induce severe perspective distortions and self-occlusions. Our analysis reveals conventional CNN-based approaches relying on single-scale receptive fields fail to capture hierarchical frequency features, while standard self-attention mechanisms tend to over-smooth critical high-frequency edge patterns. To tackle this challenge, we propose Cross-Vertical-View Network(CVVNet), a frequency aggregation architecture specifically designed for robust cross-vertical-view gait recognition. CVVNet employs a Multi-Scale Attention Gated Aggregation (MSAGA) module which consists of a High-Low Frequency Extraction module (HLFE) and a Dynamic Gated Aggregation (DGA) mechanism. Specifically, HLFE adopts parallel multi-scale convolution/max-pooling path and self-attention path as high- and low-frequency mixers for effective multi-frequency feature extraction from input silhouettes, while DGA is introduced to adaptively adjust the fusion ratio of frequency and spatial features. The integration of HLFE and DGA enables CVVNet to effectively handle distortions from view changes, significantly improving the recognition robustness across different vertical views. Experimental results show that our CVVNet achieves state-of-the-art performance, with 8.6% improvement on DroneGait and 2% on Gait3D compared with the best existing methods.
@@ 1 Introduction
In real-world surveillance scenarios, cameras are generally mounted at elevated positions [18], creating complex viewing conditions. As illustrated in Fig. 1, during normal walking sequences, subjects experience continuous distance variations (far-to-near) and vertical angle change (low-to-high) with respect to the camera. This dynamic perspective induces significant deformation in critical anatomical components (e.g., foot trajectories and joint angles), while key body regions (e.g., torso and feet) suffer from severe self-occlusion under increased vertical angles. Our experimental results demonstrate that under these challenging conditions, both skeleton-based and local texture-based feature extraction approaches exhibit substantially degraded performance. This perspective sensitivity constitutes a fundamental limitation that severely constrains the practical applicability of existing algorithms, presenting significant challenges to recognition robustness in pervasive cross-vertical viewing-angle scenarios in real-world applications.
@@
Existing methods [4, 5, 33, 6] encounter two fundamental limitations under cross-vertical-view scenarios. First, convolutional networks with fixed receptive fields predominantly focus on single-scale details, resulting in insufficient capacity for modeling hierarchical frequency information. This shortage makes it challenging to maintain spatial consistency of contours under dynamic viewpoint changes, ultimately amplifying self-occlusion effects. Second, although self-attention mechanisms offer superior low-pass filtering characteristics that effectively capture low-frequency global feature, it suffers from an inherent constraint on preserving high-frequency motion details [37, 38]. Such fragmented frequency representation creates a critical dilemma under continuous vertical viewpoint variations: convolutional architectures fail to compensate for cross-view deformations due to their weak sensitivity to global structural information, whereas self-attention mechanisms sacrifice fine-grained motion cues that are essential for discriminative identity features. The combined effect of these limitations substantially undermines the recognition robustness in complex outdoor surveillance environments, where vertical angle variations are prevalent and unavoidable.
@@
To address the aforementioned challenges, we propose a Cross-Vertical-View Network (CVVNet) to address the generalization bottleneck of existing methods in cross-view scenarios by constructing dual mechanisms of multi-frequency feature separation and spatially adaptive gated fusion. The key component of CVVNet is the Multi-Scale Attention Gated Aggregation (MSAGA) module which consists of a High-Low-Frequency feature extraction (HLFE) module and a Dynamic Gated Aggregation (DGA) mechanism. In our HLFE module, the low-frequency branch employs spatial compression attention to disentangle the global structural components of feature contours, while the high-frequency branch utilizes depthwise separable convolution to extract local dynamic components. Furthermore, we use DGA to adaptively adjust the fusion ratio of frequency and spatial features.
Our main contributions are summarized as follows.
@@
  • We design a CVVNet to achieve high-low frequency feature representations for both cross-vertical-view and outdoor recognition scenarios.
@@
  • We propose MSAGA to extract high-low frequency information from gait features and dynamically modulate the fusion ratios of different frequency components.
@@ 3.1 Overview
The framework of the proposed method is illustrated in Fig. 2. We first employ a 3D convolutional layer to extract shallow features. Subsequently, MSAGA is utilized to extract and aggregate both high- and low-frequency information from gait features. The processed features are then further refined through P3D blocks [8] to enhance their representations. Following this, a temporal max pooling operation is applied to aggregate the feature map sequence across the temporal dimension, outputting a holistic representation of the input gait sequence.
@@ 3.2 Frequency Feature Extraction based on HLFE
As shown in Fig. 4, HLFE consists of two branches for high frequency and low frequency feature extraction.
@@ 4.1 Datasets and Implementation Details
Implementation Details. Our network implementation and reproduction experiments of state-of-the-art models are both deployed through the OpenGait codebase [7], and are conducted on a system with four NVIDIA RTX 3090 GPUs. We first integrate the DroneGait dataset into this framework, following the preprocessing methodology described in [4]. The input silhouettes are aligned and resized to $64\times 44$ pixels, and then organized into sequences of length $30$ as model inputs. Both the loss functions and data augmentation strategies of CVVNet adhere to the configurations established in GaitBase [7]. To ensure fair training, we set the batch size for all models to 8×16 for DroneGait and 32×4 for Gait3D, while keeping other settings consistent with the original reproduced models.
@@ 4.2 Results and Analysis
Evaluation on Gait3D. The results on the Gait3D dataset further validate our hypothesis: enhancing viewpoint generalization directly translates to improved accuracy in complex outdoor environments. As shown in Tab 3, CVVNet achieves state-of-the-art performance across all metrics, attaining 75.6% Rank-1, 89.1% Rank-5, and 68.3% mAP. This represents a 1.2% Rank-1 improvement over DeepGaitV2 (74.4%) and a 2.5% mAP gain, demonstrating particularly strong advantages in handling real-world variations like occlusions and viewpoint shifts.
@@
Cross-domain Evaluation. Our bidirectional cross-dataset evaluation between Gait3D and DroneGait validates CVVNet’s superior domain generalization. As show in Tab 2 and Tab 4, when trained on DroneGait-Low and tested on Gait3D, CVVNet achieves a state-of-the-art 20.9% Rank-1 accuracy, surpassing DeepGaitV2’s previous best result (20.5%) by 0.4% while attaining 14.2% mAP—a 0.3% improvement over DeepGaitV2. Notably, it outperforms QAGait by 5.4% absolute Rank-1 accuracy, confirming effective transfer of drone-learned viewpoint invariance to outdoor scenes. In the reverse evaluation trained on Gait3D and tested on DroneGait-High, CVVNet demonstrates consistent superiority: 45.9% NM, 48.4% BG, and 47.5% CL accuracy under high-vertical-view conditions, exceeding DeepGaitv2 by 0.7-0.8% across metrics. For low-vertical-view scenarios, it reaches 98.4% NM and 99.5% BG accuracy, outperforming GaitBase’s prior records by 1.2%. This bidirectional dominance stems from our dual-branch architecture’s adaptive feature disentanglement, effectively bridging drone-ground domain gaps.
@@ 4.4 Visualization
We draw the heatmap to highlight the viewpoint-specific feature activation in Fig. 5. CVVNet dynamically prioritizes lower-body regions at low vertical angles and shifts attention to upper-body zones at high angles. Specifically, in low vertical views, the network focuses on dynamic details of the legs and feet, capturing crucial gait patterns that distinguish individual walking styles. In contrast, under high vertical views, the attention shifts to the head and torso, where distinctive features are more pronounced. This ability to adaptively adjust feature extraction priorities ensures that CVVNet effectively integrates multi-scale features, enhancing its robustness across varying perspectives and its generalization in complex environments. The spectral analysis is shown in Fig. 6, which reveals distinct frequency characteristics: conventional convolution-based methods (GaitGL, GaitBase, DeepGaitV2) exhibit limited frequency diversity with concentrated central energy. In contrast, our CVVNet’s MSAGA architecture demonstrates superior multi-scale frequency extraction, capturing both low- and high-frequency components. Increasing MSAGA layers progressively enriches the spectral content, enabling more comprehensive gait representations.
Figure 6: Frequency Spectrum Analysis of Different Gait Feature Extraction Methods. CVVNet-MSAGA1,3,5 indicates the module is in the 1st, 3rd, and 5th layers of Stage 2 in the model.
@@ 5 Conclusion
6 Acknowledgments
This work was supported in part by the National Natural Science Foundation of China (No. 62371142, Grant No. 62273107) and Guangdong Basic and Applied Basic Research Foundation (No. 2024A1515010404).